What is the number of moles of O-atom in 126 amu of HN03 (Nitric acid)?
1
Expert's answer
2020-11-12T06:36:00-0500

n(HNO3)=123/72=1.71

%(O)=48/72=0.67

0.67*1.71=1.14 moles

Answer: 1.14 mole